OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 4) and whose slope is 1/2

OpenStudy (anonymous):

y = mx+b where m = slope so y = (1/2)x + b plug in the given values of x and y to get 4 = (1/2)(2)+b solve for b

OpenStudy (anonymous):

would it be y – 2 = 4(x – 1/2) ??

OpenStudy (anonymous):

no because if you multiplied that out you would get \[y-2=4x-2\]\[y = 4x\] Which means that it has a slope of 4, not 1/2

OpenStudy (anonymous):

oh so its y – 4 = 1/2(x – 2) (or + 2, im not sure!)

OpenStudy (anonymous):

yes. \[y-4=.5(x-2) = .5x-1\]\[y=.5x+3\]Check your answer by plugging in\[(4)=.5(2)+3 = 1+3 = 4\]